Caleb has a metal stand that holds 600 heavy boxes. He can put 60 boxes on each shelf. If 60​% of the metal stand is​ empty, how many shelves are filled with​ boxes?
1
Expert's answer
2020-05-03T14:40:26-0400

Not empty: (100%-60%) * 600 = 0.4 * 600 = 240.

Number of shelves: 240/60= 4

Answer: 4